When an electric current is passed through acidulated water 112ml of hydrogen gas at step collects at the cathode in 965sec. The current passed in ampere is?

Grade:12

4 Answers

Khimraj
5 years ago
mole of hydrogen = 112/22400 = 1/200 = 0.005
charge required = 0.005*96500  = 482.5 C
THEN
I*t = 482.5
I*965 = 482.5
So I = 0.5A
Hope it clears. If you like answer then please approve it.
ANISHA GHOSH
5 years ago
2H2O+ 2e-  → H2 + 2OH-
no. of moles=112/22400=1/200 mol 
Q=(1/200)*96500*2         (As 2 electrons are involved)
we know,
It=Q
=> I*965 = 965
=> I = 1A
hope i cleared ur doubt...
 
Badarunissa
5 years ago
The first answer is wrong. For 1mole hydrogen 2F is needed...that is 2×96500
So .005 mole requires .005×2×96500C
=965C
Q=It
I=Q÷t
 =965÷965=1Ampere
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached answer to your problem below.
 
2H2O+ 2e-  → H2 + 2OH-
no. of moles=112/22400=1/200 mol
Q=(1/200)*96500*2         (As 2 electrons are involved)
we know,
I = Q/t
=> I = 965/965
=> I = 1A
 
Hope it helps.
Thanks and regards,
Kushagra